Perform the indicated operation. Write the result in simplest form.15 3/5÷4 4/5
Answer by stanbon(57203)

Perform the indicated operation. Write the result in simplest form.
15 3/5÷4 4/5
-----
(5*15+3)/5 / (5*4+4)/5
---------
Invert the denominator and multiply to get:
(5*15+3)/5 * 5/(5*4+4)
-----
(78/5)*(5/24)
-----
= 78/24
---
= 39/12
---
= 13/4
============
Cheers,
Stan H.
===============